SOHO DAILY REPORT Operations Day - 198/2003 (17 July 03) Mission Day 2785 SOHO FLIGHT OPERATIONS TEAM, NASA/GSFC Ref. #: SH-FOT-DR-030717 ==================================================================== PASS #1 Canberra DSS46 0113-0710 PASS #2 Madrid DSS66 0650-1401 PASS #3 Goldstone DSS27 1408-2315 ********************************************************************* PASS #1 Canberra DSS46 0113-0710 Anomalies: None NRT: None Activities Summary: - SSR Dump - Collected MDI-M (0355-0411) - Collected MDI-H (0411-0631, 0638-0710) - OBT Distribution to SWAN, CELIAS - Collected Ranging Data - SSR SEF Reset (x2) ------------------ PASS #2 Madrid DSS66 0650-1401 Anomalies: G03-0490 0700 D66 unable to maintain lock on spacecraft receiver #2 due to problems with the site's transmitter. Lost a total of 39 minutes of command ability. DR #M101737. G03-0491 1228 History files generated on String #4 from 198/0400 until 1513 corrupted due to a failure of the front end. Rebooted the front end, and the entire string, to correct problem. Operations moved over to String #5. NRT: 1238-1321 Activities Summary: - Collected MDI-M (1201-1340) - Collected MDI-H (0650-1201) - SSR SEF Reset - SSR Record ------------------ PASS #3 Goldstone DSS27 1408-2315 Anomalies: None NRT: 1933-1943 2114-2213 Activities Summary: - SSR Dump - Collected MDI-M (1434-2256) - CRS Drift Estimation Update - Uplinked delayed load CEL030717301 - New Tracking Star Slot 5 / Magnitude 6.75 - Reaction Wheel Speed Monitoring Daily Adjustment - CAE-A in the AOCS HW Config Word - HGA Table Upload - Hx Limit Update - SSR SEF Reset (x2) - SSR Record ------------------ -- End of Report --